

********************************************************************
UARS WEEKLY STATUS REPORT - 04 MAR 2005
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 57% of
each day's data amounting to nearly 577 minutes of solar observing time
per day.
***********
SUSIM
***********
During this past week, SUSIM successfully gathered daily mid (1.1 nm)
resolution, MgII, occultation, solar pointing offset, and continuous
monitoring scans for time periods with UARS telemetry or tape
recording.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 19 February 2005 (Orbit 73572, GMT Day 050) through
25 February 2005 (Orbit 73672, GMT Day 056).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 056/0510 73664 POWER OFF
The beta angle ranged between the angles of 27.6 degrees to 49.0
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 63.0 degrees on 02 March 2005 (DOY 061).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 3.91 to 3.31 degrees C. Additional
battery performance data for each battery on 18 February 2005 and 25
February 2005 are provided in Appendix D.
The clock error ranged from 10.2 to -10.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to monitor the batteries closely. The MPS
continues to utilize the VT-7 charge rate.
The FOT has completed their CMS software acceptance testing. All
tests results were positive and the official release of the new
software was installed on the backup CMS. With the install complete,
the FOT will run a few simulated days of operations to confirm the
software. If all goes well, then the new release will be put onto the
operational CMS. The day-in-the-life simulations is awaiting
simulator support, which should be available next Wednesday.
MEETINGS
The FOT attended the second clock rollover meeting on 22 February 2005
to discuss the activities required for the UARS flight software epoch
year change. This meeting will be held biweekly.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 March 2005 (DOY 080). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 25 February 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
73641 054/1447 05-050 OBC PMON ACTION CLEAR
73652 055/0902 05-051 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
73664 056/0510 05-052 HALOE POWER OFF (HI-BETA)
73674 056/1947 05-053 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 25 February 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 25 February 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
73652 055/0927 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
73674 056/1950 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 25 February 2005
18 February 2005 - GMT Day 049 Beta = 22.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.11/-3.05 +00.0/+00.0 07.0 0.00 0.0
2 235.2/ 67.2 +10.06/+9.06 +28.8/-12.8 26.7 1.07 19.9
3 -16.8/-235.2 +6.44/+5.79 +25.6/-12.0 26.7 1.05 18.4
25 February 2005 - GMT Day 056 Beta = 49.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.79/-3.05 +00.0/+00.0 07.0 0.00 0.0
2 168.0/ 56.0 +10.40/+9.06 +26.4/-15.2 26.7 1.11 17.0
3 11.2/- 89.6 +7.09/+5.15 +22.4/-14.4 26.7 1.07 16.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 25 February 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-005 FOT 73587 051/0038 30832 Generic Late Acquisition
05-005 FOT 73591 051/0730 30832 Generic Late Acquisition
05-010 FOT 73616 052/2218 38591 Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-005 FOT See List Feb 05 30832 Generic Late Acquisition
05-002 MPS 73008 012/1251 N/A BATT 3 Diff Voltage High
05-010 FOT 73616 052/2218 38591 Late Acquisition